### v1

# MobileVLM: A Fast, Strong and Open Vision Language Assistant for Mobile Devices

Abstract
We present MobileVLM, a competent multimodal vision language model (MMVLM) targeted to run on mobile devices. It is an amalgamation of a myriad of architectural designs and techniques that are mobile-oriented, which comprises a set of language models at the scale of 1.4B and 2.7B parameters, trained from scratch, a multimodal vision model that is pre-trained in the CLIP fashion, cross-modality interaction via an efficient projector. We evaluate MobileVLM on several typical VLM benchmarks. Our models demonstrate on par performance compared with a few much larger models. More importantly, we measure the inference speed on both a Qualcomm Snapdragon 888 CPU and an NVIDIA Jeston Orin GPU, and we obtain state-of-the-art performance of 21.5 tokens and 65.3 tokens per second, respectively. Our models are available at: https://github.com/Meituan-AutoML/MobileVLM
1 Introduction
Large multimodal models (LMMs), especially the family of Visual Language Models (VLMs), rise as a promising research direction for building general-purpose assistants due to their substantially enhanced capability in both perception and reasoning [127]. However, it has been challenging to connect the representations of the pre-trained large language models (LLMs) [49, 131, 26, 118, 115] and the vision models to unveil the cross-modality properties, such as visual question answering, image captioning, visual knowledge reasoning and conversation, etc.
Remarkable performance on this task can be witnessed in GPT-4V [90] and Gemini [45], and the evaluation of their abilities can be found in [1, 41]. However, very limited technical details are available for these proprietary models. Simultaneously in the research community, a line of language tuning methods have been proposed [2, 66, 30, 76, 74, 78, 5, 133, 126, 122]. For instance, Flamingo [2] exploits visual tokens to condition the frozen language model via gated cross-attention layers. BLIP-2 [66] argues that such interaction is insufficient and introduces a lightweight querying transformer (called Q-Former) that extracts the most useful features from the frozen vision encoder and feeds them directly into the frozen LLM. MiniGPT-4 [133] aligns a frozen visual encoder from BLIP-2 [66] with a frozen language model Vicuna [118] via only one projection layer. Independently, LLaVA [76] applies a simple trainable projector that converts the vision features into embedding tokens, which have the same dimension as the word embeddings to be processed by the language model altogether.
Noticeably, training strategies also exhibit a shift to accommodate the large-scale multimodal data of great diversity. LLaVA may be the first attempt to replicate the instruction tuning paradigm from LLMs to the multimodal scenario. To generate multimodal instruction-following data, it feeds textual information such as captions and bounding-box coordinates of images to language-only GPT-4 [89]. MiniGPT-4 [133] is firstly trained on a combined image captioning dataset and then fine-tuned on a curated alignment dataset of image-text pairs. InstructBLIP [30] enforces vision-language instruction tuning based on the pre-trained BLIP-2 model, where the Q-Former is trained on a diverse set of datasets organized in an instruction-tuning format. mPLUG-Owl [126] introduces a two-stage training strategy where the visual part is pre-trained first and the large language model LLaMA [115] is then fine-tuned with LoRA [51] with instruction data from various sources.
Despite the advances mentioned above of VLMs, there is a natural demand to enable cross-modality capacities in resource-constrained scenarios. Gemini [45] surpasses state-of-the-art performance on a range of multimodal benchmarks and introduces mobile-scale VLMs with 1.8B and 3.25B parameters for low-memory devices. Common compression techniques such as distillation and quantization are also exploited for this purpose. We aim to build the first open, mobile-scale VLMs trained using public datasets and available techniques to achieve visual perception and reasoning, customized for resource-constrained platforms. Our contributions are as follows:
1.
We present MobileVLM, a full-stack remake of multimodal visual language models tailored for mobile scenarios. To our knowledge, we are the first to provide a detailed, reproducible, and strong vision language model from scratch. With controlled and open-source datasets, we build a set of high-performance foundation language models and multimodal models.
2.
We make extensive ablation studies on the design of visual encoders and systematically evaluate the VLM performance sensitivity on various training paradigms, input resolutions, and model sizes.
3.
We design an efficient projector between visual and text features, which better aligns multimodal features while reducing the inference budget.
4.
Our model is crafted to run efficiently on mobile, low-power devices, with a measured speed of 21.5 tokens/s on a Qualcomm mobile CPU and 65.3 tokens/s on a Jeston Orin GPU respectively.
5.
Our models perform comparably on a large body of VLM benchmarks, attesting their potential in numerous tasks in practice. Although we mainly focus on edge scenarios, our model outperforms many recent VLMs, which can only be supported by powerful GPUs in the cloud.
2 Related Work
2.1 Vision Transformer
Vision Transformer [34] is now the dominant backbone for visual perception. Its sequel methods such as Swin [81], DeiT [114], PVT [121], and Twins [22] have upgraded its original architectural design to strengthen its representational power and efficiency. The pre-training paradigm has also experienced several shifts, from image-supervised learning (i.e. labeled classification) to unsupervised learning like masked auto-encoders [47], and most recently to language-supervised training as advertised by CLIP [97], which empowered ViTs with unprecedented zero-shot capability. VLMo [6] enhances CLIP with unified multimodality training.
2.2 LLMs
Large language models often come with billions of parameters and are pre-trained on extremely extensive text corpora, exhibiting emergent capabilities [123] that have not been witnessed before. They have reshaped the field of natural language processing and are being used in a wide range of applications. To date, proprietary LLMs like GPT-4 [89] prevail over open-sourced models. Nevertheless, the community is exuberant with the continuous model releases, including GLM [35], BLOOM [65], OPT [131] and LLaMA series [115, 116]. Many recent works [132, 4] have been built on top of them.
Noticeably, there is a trend to build smaller language models, i.e., whose parameters are around 1B or fewer. To name a few, GPT-Neo [9], Pythia [7], GALACTICA[112], OpenLLaMA [43], Phi [46, 70], Qwen [4] all ship language models at such sizes. Although privately collected high-quality data can significantly boost the performance of LLMs [70, 4], our target is to build reproducible and efficient models, hence we do not utilize any non-public data for our research.
2.3 VLMs
| Model | Vision Encoder | Language Model | Cross-modality Design | Multimodal Training Corpora |
| CLIP [97] | ViT, ResNet | Transformer | Linear-Projection | WebImageText [97] (400M image-text pairs) |
| BLIP [67] | ViT | MED∗ | Cross-Attention | COCO [73], VG [62], CC3M [103], CC12M [13], LAION [102] |
| BLIP-2 [66] | CLIP/Eva-CLIP ViT | OPT, Flan-T5 | Q-Former | same as BLIP |
| InstructBLIP [30] | ViT-G/14 | Flan-T5, Vicuna | Q-Former w/ FC | 13 held-in out of 26 public datasets |
| Flamingo [2] | NFNet-F6 [59] | Chinchilla | Perceiver-Resampler [57] | M3W(43M), ALIGN(1.4B) [58], LTIP (312M), VTP (27M) |
| LLaVA [76] | CLIP ViT-L/14 | Vicuna 7B/13B | Linear-Pprojection | filtered CC-595K from CC3M [103], LLaVA-Instruct-158K |
| LLaVA-1.5 [74] | CLIP ViT-L@336 | Vicuna-7B/13B | MLP | a subset of InstructBLIP (1.2M) |
| MiniGPT-4 [133] | EVA-CLIP ViT-G/14 | Vicuna-7B | Q-Former | LAION, CC, SBU [91] |
| Shikra [15] | CLIP ViT-L/14 | Vicuna-7/13B | FC-layer | Flickr30K [95], RefCOCO [61], VG, Visual-7W [85] |
| mPLUG-Owl [126] | CLIP ViT-L/14 | LLaMA-7B | Cross-Attention | LAION-400M, COYO-700M [11], CC, COCO |
| KOSMOS-2 [94] | VLMo [6] | MAGNETO [119] | Perceiver-Resampler | GRIT (curated with LAION-2B, COYO-700M) |
| QWen-VL [5] | Openclip ViT-bigG [54] | Qwen-LM | Cross-Attention | multi-tasking datasets (Captioning, VQA, Grounding, etc. ) |
| ShareGPT4V [16] | CLIP ViT-L/14@336 | Vicuna-7B | MLP | ShareGPT4V (100K by GPT-4V, 1.2M by its learned model) |
| MobileVLM (ours) | CLIP ViT-L/14@336 | MobileLLaMA | LDP | same as LLaVA-1.5 [74] |
Throughout recent years, a school of vision language models has rapidly emerged. Table 1 summarizes them in a detailed comparison regarding architectures, cross-modality design, and training corpora.
Architecture choices. As a consequence of the intimidating training cost of large language models, most language models used in VLMs are pre-trained open-source models like OPT [131], Flan-T5 [26], Chinchilla [63], Vicuna [118] and LLaMA [115]. QWen adapts LLaMA with custom variations [4] to obtain an LLM on their own.
Visual backbones in VLMs are typically vision transformer [34], but pre-trained in various strategies [97, 66, 37]. Most VLMs prefer CLIP-fashioned ViT [97] trained with natural language supervision. Flamingo picks NFNet-F6 [59]. KOSMOS chooses VLMo [6] instead.
Dataset-centric-ness. The construction of training data has become increasingly crucial. It is quite common to utilize millions of text-image pairs in the line of VLMs, where the new datasets are usually released alongside their corresponding new models. To name a few, apart from an enhanced visual receptor and novel language model called Qwen-LM [4], the multilingual multimodal Qwen-VL [5] additionally aligns the image with caption and box tuples, which sets a new record of generalist models. PALI [19] and PALI-X [18] consume an internal multi-language image-text dataset called WebLI at a scale of 12 billion. Most recently, observing the constraints of current image-text datasets like hallucination and inaccurate descriptions, ShareGPT4V [16] exploits GPT-4V [90] for generating 1.2M high-quality image-text pairs with which can surpass the LLaVA series. Similarly built with GPT-4V, LVIS-INSTRUCT4V [120] helps LLaVA-1.5 to gain substantial improvement on various VLM benchmarks.
2.4 Model Compression for LLMs
Large Language Models (LLMs) have brought a paradigm shift in natural language processing, while their colossal size and computational requirements pose significant challenges for real-world deployment, particularly in environments with limited resources. The size of these models often results in high memory usage and slow processing speeds. Additionally, the energy requirements for training and operating these models raise sustainability concerns. These challenges are becoming more pressing as LLMs continue to grow in size and complexity. In response to these challenges, model compression has emerged as a crucial research area, which aims to reduce the size and computational demands of models without significantly compromising their performance. These techniques include but not limited to model pruning [84, 38, 109], quantization [39, 124, 68], knowledge distillation [130] and low-rank decomposition [125].
2.5 VLM Benchmarks
Systematic and comprehensive evaluations of different VLMs are of great necessity. To this end, POPE [71] provides a benchmark for evaluating hallucination in VLMs, which formulates the evaluation as a binary classification task that prompts the VLMs to answer whether the object exists. In contrast, GQA [52] mainly centers around the VLMs’ abilities in real-world reasoning, scene understanding, and compositional question answering. TextVQA [105] consists of questions related to text in the image, evaluating the OCR and reasoning abilities of models. ScienceQA [83] consists of multimodal multiple-choice questions covering scientific topics, e.g., natural science, social science, and language science, which requires VLMs to integrate common sense into reasoning. MME [40] measures both the perception and cognition abilities of VLMs, it covers a total of 14 subtasks varying from coarse-grained to fine-grained ones. MMBench [80] is a methodically constructed multimodal dataset, which collectively covers a diverse spectrum of 20 fine-grained skills and involves a circular evaluation strategy with the incorporation of ChatGPT [88].
2.6 Embodied AI
Our work is closely related to Embodied Artificial Intelligence. Being one of the central goals of Artificial General Intelligence, embodied AI strives to build egocentric intelligence systems that can interact with their surroundings with perception, reasoning, and planning capabilities [36]. Recently, the emerging large vision language models [86, 106, 108] allow embodied AI agents to resolve the relevant tasks like embodied question answering [32], and vision-language navigation [69, 92] in a highly end-to-end fashion.
3 MobileVLM
3.1 Overall Architectural Design
With the primary goal of achieving efficient visual perception and reasoning for resource-limited edge devices in mind, we design the overall architecture of MobileVLM as illustrated in Figure 1. It contains three components: 1) a visual encoder, 2) a tailored LLM for edge devices (MobileLLaMA), and 3) an efficient projector (termed “lightweight downsample projector”, LDP) that aligns the visual and the textual space.
Taking an image as input, the vision encoder extracts the visual embeddings for image perception, where denotes the number of image patches and denotes the hidden size of visual embeddings. To alleviate the efficiency issues arising from prolonged image tokens, we scheme a lightweight projector for visual feature compression and visual-text modal alignment. It converts into the word embedding space with an appropriate input dimension of the subsequent language model as below,
Thus we obtain the image tokens and text tokens , where denotes the number of text tokens and denotes the hidden size of the word embedding space. Observing that LLM occupies the most computation and memory consumption in the current design paradigm of MLLMs, we tailor a series of inference-friendly LLMs that enjoy advantages in speed for mobile applications. It predicts the response conditioned on the multimodal input in an autoregressive manner, where denotes the output token length. This process can be formulated as,
3.2 Visual Encoder
Based on empirical analysis later shown in Sec 5.1, we leverage the pre-trained CLIP ViT-L/14 [97] with a resolution of 336336 as our visual encoder . The Vision Transformer (ViT) [34] dissects images into uniform-sized patches, applies a linear embedding to each, and integrates positional encodings before feeding the resultant vector sequence into a canonical Transformer encoder. Typically, a classification token is appended to this sequence for subsequent categorization tasks.
3.3 MobileLLaMA
For the language model, we downscale LLaMA [115] to facilitate the off-the-shelf deployment, i.e., our models can be seamlessly supported by almost all popular inference frameworks. Moreover, we evaluate the model latency on the edge devices to guide the model design. Neural architecture search [77, 24, 50, 25, 12] would be a better choice, but for the time being we keep it as our future work. The detailed setting of our architecture is shown in Table 2.
| Model | Blocks | Dim | Heads | Context length |
| MobileLLaMA 1.4B | 24 | 2048 | 16 | 2k |
| MobileLLaMA 2.7B | 32 | 2560 | 32 | 2k |
Specifically, we utilize the sentence piece tokenizer [63] from LLaMA2 [116] with a vocabulary size of 32000 and train the embeddings from scratch. This is beneficial for performing future distillation without further pain. The context length used at the pre-train stage is 2k for all models due to limited resources. However, the context window can be further scaled to 8k for inference, as indicated by [17]. The detailed settings of other components are listed below.
3.4 Efficient Projector
The projector between the vision encoder and the language model is critical in aligning multimodal features. There are two existing paradigms: Q-Former [133, 66] and MLP projection [76, 74]. Q-Former explicitly controls the number of visual tokens per query to force extracting the most relevant visual information. However, it inevitably loses the spatial positional information of tokens and suffers from slow convergence. In addition, it is inefficient for the inference on edge devices. In contrast, MLP retains the spatial information but it usually includes useless tokens such as the background. For an image of with a patch size of , there are visual tokens to be injected into the LLM model, which greatly slows down the overall inference speed. Inspired by [23], we can utilize convolutions to enhance the positional information and encourage local interaction of the vision encoder. Specifically, we explore mobile-friendly operations based on depth-wise convolution (the simplest form of PEG [23]), which is efficient and well-supported by various edge devices.
To keep spatial information and to minimize the computational cost, we make use of a convolution with a stride of 2, which reduces 75% visual tokens. This design significantly boosts the overall inference speed. However, our experimental result indicates that solely downsampling the tokens severely deteriorates the performance on downstream tasks, such as OCR. To alleviate this effect, we utilize a more powerful network instead of a single PEG. The detailed architecture of the efficient projector, called Lightweight Downsample Projector (LDP), is illustrated in Figure 2. Note that, this projector network only contains less than 20M parameters and runs about faster than the visual encoder.
We use Layer Normalization instead of Batch Normalization [56] to make training stable and not affected by the batch size. Since the projector is already very light-weight, therefore, we don’t adopt recent mobile reparameterization designs [117, 21].
Formally, LDP (denoted as ) takes the visual embeddings as input, and outputs the efficiently extracted and aligned visual tokens as,
| Model | Common Sense Reasoning | Language Understanding | ||||||||
| ARC(c/e) | BoolQ | RTE | Winogrande | TruthfulQA | HellaSwag | PIQA | MMLU | Avg. | ||
| INCITE 3B (V1) [29] | 0.32 / 0.68 | 0.67 | 0.52 | 0.63 | 0.21 / 0.33 | 0.48 | 0.74 | 0.2675 | 0.4848 | |
| OpenLLaMA 3B (V1) [43] | 0.34 / 0.69 | 0.68 | 0.58 | 0.62 | 0.22 / 0.35 | 0.49 | 0.75 | 0.2695 | 0.4990 | |
| MobileLLaMA 2.7B | 0.32 / 0.68 | 0.61 | 0.59 | 0.63 | 0.23 / 0.36 | 0.48 | 0.75 | 0.2730 | 0.4923 | |
| TinyLLaMA 1.1B (2T) [93] | 0.23 / 0.57 | 0.59 | 0.55 | 0.57 | 0.23 / 0.39 | 0.40 | 0.70 | 0.2541 | 0.4484 | |
| Galactica 1.3B [112] | 0.28 / 0.62 | 0.62 | 0.52 | 0.55 | 0.25 / 0.41 | 0.34 | 0.63 | 0.2675 | 0.4488 | |
| OPT 1.3B [131] | 0.24 / 0.57 | 0.56 | 0.51 | 0.59 | 0.24 / 0.39 | 0.41 | 0.71 | 0.2461 | 0.4466 | |
| Pythia 1.4B [7] | 0.26 / 0.61 | 0.63 | 0.52 | 0.57 | 0.23 / 0.39 | 0.40 | 0.71 | 0.2568 | 0.4577 | |
| MobileLLaMA 1.4B | 0.26 / 0.61 | 0.53 | 0.53 | 0.59 | 0.21 / 0.35 | 0.43 | 0.73 | 0.2497 | 0.4490 | |
4 Experiment
4.1 Training
The whole reproducible training process is composed of three stages. Firstly, we pretrain LLM foundation models on the text-only dataset RedPajama v1 [29]. Secondly, we perform supervised fine-tuning (SFT) following Vicuna [20] on a dataset of multi-turn dialogues between humans and ChatGPT from third-party platforms. Lastly, we train our vision large models using multimodal datasets.
Language model pre-training. Since our target is training deployable models for inference, we do not strictly follow the efficient combination of model capacity and tokens from scaling law [60]. To make our work reproducible, all the models are trained on 1.3T tokens111Applying the tokenizer of [116] on this dataset generates about 1.3T tokens. from the RedPajama v1 [29] dataset only. This benefits further research by enabling controlled experiments. We apply the same sampling ratio of different datasets as [115], which is shown in Table 12 (see Appendix). The common autoregressive loss is adopted. We utilize a global batch size of 5,242,880.
The peak learning rate is set to 3 10-4 and it decays to 3 10-5 following the cosine strategy. We warm up with 2000 iterations. We use the AdamW optimizer [82] with and and a weight decay regularization value of 0.1. The gradient clipping threshold is set to 1.0. We adopt the Pytorch lightning framework with DeepSpeed backend [99]. Specifically, we utilize ZERO 1 [98] and gradient accumulation to achieve a training speed of 18800 TGS (Tokens per second for a GPU) for the 1.4B model and 8500 TGS for the 2.7B model, on 20 nodes equipped with 8 NVIDIA Tesla A100 GPUs each.
Furthermore, we also favor Flash Attention V2 [31] to alleviate the I/O bottleneck and to train faster. We randomly shuffle the data to disturb the sequential order with a fixed seed, which is vital since the training process can be intermittently interrupted and requires to be resumed. We first tokenize the raw data into IDs and serialize them into many bucket files. We then utilize memory mapping to deliver a desired I/O speed. Besides, we pack different sentences together, where an EOS token is inserted to set different sentences apart. Due to limited resources, we do not try the design of InternLM [113], which may further improve the model performance by disabling such packing. The overall training loss decreases as the consumed tokens increase and is shown in Figure 3.
SFT on language models. As clarified by LLaMA-2 [116], fewer higher-quality examples from their vendor-based annotation efforts significantly improve the results. We are thus motivated to finetune our MobileLLaMA on a high-quality dataset with supervised learning. Vicuna [20] fine-tunes LLaMA on user-shared conversations collected from ShareGPT, which is widely used as a language module for multimodal model building, but their training dataset is not released. We employ a dataset of multi-turn dialogues between humans and ChatGPT from third-party platforms [53] which has been cleaned through a process of format standardization and quality refinement. The SFT data is organized following the Vicuna format, where each sample consists of a dialogue including several user prompts and ChatGPT answers. As shown in Table 14 (see Appendix), a special token is utilized to separate the assistant’s answer and the next round of user prompts. For the training details, we use a cosine learning rate schedule without weight decay, a global batch size of 128, and a sequence length of 2048 tokens. We utilize an autoregressive objective and perform backpropagation only on answer tokens. To achieve better performance in downstream tasks, we conducted experiments to select the appropriate hyperparameters. We fine-tune for 3 epochs with a learning rate of for MobileLLaMA 1.4B, and 2 epochs with a learning rate of for MobileLLaMA 2.7B. The training loss decreases with iterations as shown in Figure 4. To be later shown in Sec. 5.4, our empirical performance on downstream tasks demonstrates that high-quality SFT data is essential to aligning LLMs with dialogue-style instructions.
| Method | LLM | Res. | PT | IT | GQA | SQA | VQA | POPE | MME | MMB |
| Openflamingo [3] | MPT-7B | 336 | 180M | - | – | – | 33.6 | – | – | 4.6 |
| BLIP-2 [66] | Vicuna-13B | 224 | 129M | - | 41.0 | 61.0 | 42.5 | 85.3 | 1293.8 | – |
| MiniGPT-4 [133] | Vicuna-7B | 224 | 5M | 5K | 32.2 | – | – | – | 581.7 | 23.0 |
| InstructBLIP [30] | Vicuna-7B | 224 | 129M | 1.2M | 49.2 | 60.5 | 50.1 | – | – | 36.0 |
| InstructBLIP [30] | Vicuna-13B | 224 | 129M | 1.2M | 49.5 | 63.1 | 50.7 | 78.9 | 1212.8 | – |
| Shikra [15] | Vicuna-13B | 224 | 600K | 5.5M | – | – | – | – | – | 58.8 |
| mPLUG-Owl [126] | LLaMA-7B | 224 | 2.1M | 102K | – | – | – | – | 967.3 | 49.4 |
| IDEFICS-9B [64] | LLaMA-7B | 224 | 353M | 1M | 38.4 | – | 25.9 | – | – | 48.2 |
| IDEFICS-80B [64] | LLaMA-65B | 224 | 353M | 1M | 45.2 | – | 30.9 | – | – | 54.5 |
| Qwen-VL [5] | Qwen-7B | 448 | 1.4B | 50M | 59.3 | 67.1 | 63.8 | – | 1487.6 | 38.2 |
| MiniGPT-v2 [14] | LLaMA-7B | 448 | 23M | 1M | 60.3 | – | – | – | – | 12.2 |
| LLaVA-1.5 [74] | Vicuna-7B | 336 | 558K | 665K | 62.0 | 66.8 | 58.2 | 85.9 | 1510.7 | 64.3 |
| MobileVLM 1.7B | MobileLLaMA 1.4B | 336 | 558K | 665K | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 |
| MobileVLM 1.7B w/ LoRA | MobileLLaMA 1.4B | 336 | 558K | 665K | 57.0 | 53.1 | 42.3 | 86.0 | 1143.7 | 50.4 |
| MobileVLM 3B | MobileLLaMA 2.7B | 336 | 558K | 665K | 59.0 | 61.0 | 47.5 | 84.9 | 1288.9 | 59.6 |
| MobileVLM 3B w/ LoRA | MobileLLaMA 2.7B | 336 | 558K | 665K | 58.4 | 59.0 | 46.7 | 84.6 | 1296.4 | 57.0 |
VLM training. Similar to [76, 126], the whole training process comprises two steps: pre-training and instruction tuning. This is depicted in Figure 2. During the first step, we freeze the vision encoder and LLM, focusing on training the efficient projector only. Subsequently, we fine-tune both the projector and LLM to enhance the abilities of visual understanding and expression by refining the language model via a language modeling loss function. Following Vicuna’s hyperparameters [118], we pre-train our model on the filtered CC-595K subset [76] for 1 epoch at a learning rate of and a batch size of 256. We fine-tune it on the LLaVA-Instruct-158K dataset [76] for 1 epoch at a learning rate of and a batch size of 128. Examples of our training dataset are shown in Figure 5 (Appendix C). We choose the AdamW optimizer with no weight decay and a cosine learning rate with a warmup ratio of 3%. The training takes 5 hours with 8 NVIDIA Tesla A100 GPUs for MobileVLM 1.7B, and 8 hours for MobileVLM 3B.
4.2 Evaluation of MobileLLaMA
In Table 3, we extensively assess our models on two standard natural language benchmarks, for language understanding and common sense reasoning respectively. We apply the Language Model Evaluation Harness [42] tool for the former assessment. Experimental results show that our MobileLLaMA 1.4B is on par with the most recent open-source models such as TinyLLaMA 1.1B, Galactica 1.3B, OPT 1.3B, and Pythia 1.4B. Notably, our MobileLLaMA 1.4B outperforms TinyLLaMA 1.1B which is trained on 2T tokens, twice as many as ours. At the 3B level, our MobileLLaMA 2.7B also demonstrates competitive performance to INCITE 3B (V1) and OpenLLaMA 3B (V1), while MobileLLaMA 2.7B being about 40% faster than OpenLLaMA 3B on a Snapdragon 888 CPU as shown in Table 5.
For common sense reasoning, we report the zero-shot accuracy on five prevalent benchmarks, i.e., ARCe [10], ARCc [28], BoolQ [27], Winogrande [101], and TruthfulQA [72]. Our models demonstrate strong reasoning capabilities on these benchmarks. It’s worth noting that MobileLLaMA 1.4B and 2.7B achieve the highest performance on Winogrande at both 1B and 3B levels. This indicates that our models deliver robust common sense capabilities, rather than cunningly relying on systematic biases in the datasets.
Further, we evaluate our models on several language understanding tasks, including PIQA [8], HellaSwag [128], and MMLU [48]. We report the zero-shot accuracy on PIQA and HellaSwag, and 5-shot performance on MMLU. We can see that our MobileLLaMA outperforms other models across nearly all of the benchmarks. The superior language understanding capability makes our models more suitable for downstream tasks, particularly for instruction tuning, in-context learning, etc.
| Model | Hardware | Precision |
| OpenLLaMA 3B | Snapdragon 888 | 8-bit | 3.4 | 3093 | 7.32 | 6.58 | 63.33 | ||||||||
| Snapdragon 888 | 4-bit | 2.3 | 3604 | 8.97 | 7.14 | 58.04 | |||||||||
| MobileLLaMA 2.7B | Snapdragon 888 | 8-bit | 2.7 | 3919 | 17.59 | 9.14 | 44.85 | ||||||||
| Snapdragon 888 | 4-bit | 1.5 | 3932 | 18.10 | 14.71 | 28.30 | |||||||||
| TinyLLaMA 1.1B | Snapdragon 888 | 8-bit | 1.1 | 4215 | 39.49 | 19.75 | 20.83 | ||||||||
| Snapdragon 888 | 4-bit | 0.6 | 3887 | 44.17 | 31.54 | 13.22 | |||||||||
| MobileLLaMA 1.4B | Snapdragon 888 | 8-bit | 1.4 | 3846 | 35.46 | 17.93 | 22.81 | ||||||||
| Snapdragon 888 | 4-bit | 0.7 | 3870 | 36.20 | 28.32 | 14.76 | |||||||||
| OpenLLaMA 3B | Jetson Orin | 8-bit | 3.4 | 2382 | 80.34 | 29.97 | 13.94 | ||||||||
| Jetson Orin | 4-bit | 2.3 | 3340 | 143.25 | 32.16 | 12.83 | |||||||||
| MobileLLaMA 2.7B | Jetson Orin | 8-bit | 2.7 | 3040 | 133.41 | 33.28 | 12.46 | ||||||||
| Jetson Orin | 4-bit | 1.5 | 2647 | 130.97 | 38.99 | 10.74 | |||||||||
| TinyLLaMA 1.1B | Jetson Orin | 8-bit | 1.1 | 3007 | 279.61 | 72.30 | 5.89 | ||||||||
| Jetson Orin | 4-bit | 0.6 | 3801 | 306.76 | 78.83 | 5.38 | |||||||||
| MobileLLaMA 1.4B | Jetson Orin | 8-bit | 1.4 | 3289 | 249.56 | 60.73 | 6.96 | ||||||||
| Jetson Orin | 4-bit | 0.7 | 3738 | 253.22 | 66.79 | 6.33 |
4.3 Comparison with SOTA VLMs
We evaluate the multimodal performance following LLaVA on GQA [52], ScienceQA [83], TextVQA [105], POPE [71], and MME [40]. In addition, we also exploit MMBench [80] for a comprehensive comparison. As demonstrated in Table 4, our proposed MobileVLM, despite its reduced parameters and limited training data, achieves competitive performance. In certain instances, it even obtains superior metrics compared with the previous state-of-the-art multimodal vision language models.
In particular, on the evaluation benchmarks of GQA, POPE, and MMBench, MobileVLM demonstrates parity with or superiority over the majority of 7B or larger VLMs, which proves its exceptional capabilities in image content perception, spatial and relational reasoning, and attribute understanding. Nonetheless, a notable shortfall arises due to the absence of large training datasets, such as code and textbook, which results in a discernible performance discrepancy on tasks like ScienceQA and MME. Still, there exists potential for enhancement in the model’s proficiency by reconciling text comprehension with the nuanced recognition of fine image content on TextVQA. We list visualized demonstrations in Appendix C.
4.4 MobileVLM with LoRA
Low-Rank Adaptation (LoRA) [51] can perform on par with or better than fully fine-tuned LLMs with fewer trainable parameters. We empirically investigate this practice on MobileVLM to validate its multimodal performance. Specifically, during the VLM visual instruction tuning stage, we freeze all the LLM parameters except the LoRA matrices. The updated parameters are only 8.87% and 7.41% of the full LLM for MobileLLaMA 1.4B and MobileLLaMA 2.7B respectively. For LoRA configuration, we set as 128 and the as 256. The results are shown in Table 4. We can see that MobileVLM with LoRA achieves comparable performances to those of full finetuning on 6 benchmarks, which is consistent with [51].
| Model | LM | Hardware |
Vicuna 7B | Snapdragon | 4.70 | 8.23 | 17347 | 5.36 | 0.25 | 329.89 | ||||||||||
| OpenLLaMA 3B | Snapdragon | 2.88 | 7.98 | 27530 | 8.95 | 7.22 | 84.43 | |||||||||||
| TinyLLaMA 1B | Snapdragon | 1.18 | 7.77 | 31370 | 41.70 | 18.40 | 20.70 | |||||||||||
MobileLLaMA 2.7B | Snapdragon | 2.14 | 8.43 | 27660 | 18.36 | 12.21 | 33.10 | ||||||||||
| MobileLLaMA 1.4B | Snapdragon | 1.40 | 6.82 | 34892 | 34.93 | 21.54 | 18.51 | |||||||||||
Vicuna 7B | Jetson Orin | 4.70 | 2.89 | 9281 | 367.26 | 17.74 | 19.75 | ||||||||||
| OpenLLaMA 3B | Jetson Orin | 2.88 | 2.94 | 22270 | 474.49 | 30.66 | 12.52 | |||||||||||
| TinyLLaMA 1B | Jetson Orin | 1.18 | 2.98 | 24655 | 1253.94 | 76.63 | 5.90 | |||||||||||
MobileLLaMA 2.7B | Jetson Orin | 2.14 | 3.11 | 15678 | 440.60 | 38.34 | 8.31 | ||||||||||
| MobileLLaMA 1.4B | Jetson Orin | 1.40 | 3.32 | 17712 | 667.69 | 65.27 | 5.14 |
4.5 Latency Measurement on Mobile Devices
We evaluate the inference latency of MobileLLaMA and MobileVLM both on a Realme GT mobile phone and an NVIDIA Jetson AGX Orin platform. The mobile phone is equipped with a Snapdragon 888 SoC and 8GB RAM, which gives out 26 TOPS computing power. Orin is equipped with 32GB of memory, offering an impressive 275 TOPS. It operates on CUDA version 11.4, which supports the latest parallel computing technology for enhanced performance. We exploit llama.cpp [44] as the inference framework for both devices.
[…]
According to the measurement on the Snapdragon 888 SoC shown in Table 5, our MobileLLaMA 2.7B significantly outperforms OpenLLaMA 3B at the same quantization precision. Notably, using two-thirds consumption of RAM and half the loading time, MobileLLaMA 2.7B achieves twice the inference speed of OpenLLaMA 3B. For the smaller model, although the quantized MobileLLaMA 1.4B is 23% larger than that of TinyLLaMA 1B, its inference speed is only 11% slower than TinyLLaMA 1B in terms of both Eval speed and Total inference time, which demonstrates that our model benefits from better architectural design. On the Orin platform, we can draw the same conclusion that MobileLLaMA is quite competitive to the open-source language models at similar parameter scales.
MobileVLM Due to the limitation of llama.cpp , VLMs are split into the vision encoder and the language model, which are loaded separately during the inference stage. For the inference precision, the language model is quantized to 4-bit, while the vision encoder and the projector keep the original precision. The prompts of VLMs are composed of textual prompts and visual prompts. For a fair comparison, LLaVA-v1.5 and MobileVLMs take the picture with the same resolution of as the visual prompt, along with the sentence “What is in the picture?” as the textual prompt. Meanwhile, the patch sizes of their vision encoder are fixed to 14. For LLaVA-v1.5, the number of input tokens (denoted as ) for the language model is composed of visual tokens and some textual tokens. For MobileVLMs, the number of visual tokens has dropped to due to the design of an efficient projector as in Sec. 3.4. For all the tested VLMs, the number of output tokens (denoted as ) is set to 256. The total inference time of VLMs is calculated specifically as,
where , , and are measured in tokens per second. is the time it takes to “tokenize” (sample) the prompt message. denotes the time LMs take to handle the processed tokens. is the time needed to generate all response tokens, measured only when the LM starts emitting tokens. and refer to the time of loading model and the entire time consumed by a single inference respectively. Other overheads are mostly negligible.
The inference latency comparison of various VLMs is shown in Table 6. It can be found that of MobileVLM-336 is slightly slower than MobileLLaMA in Table 5 since extra vision tokens increase the time consumption and consequently slow down the generation of output tokens.
Most importantly, we are surprised to find of MobileVLM equipped with MobileLLaMa 2.7B and MobileLLaMa 1.4B outperform LLaVA-v1.5 consistently on Snapdragon SoC and Jetson Orin. It is worth noting that although of MobileVLM (MobileLLaMa 1.4B) is slower than that of LLaVA-v1.5 (OpenLLaMA 1B) on Orin, it achieves a faster inference time. The main reason is that of MobileVLM is only a quarter of LLaVA-v1.5 in Equation 4 while other variables are comparable. This precisely proves the importance of efficient projection that reduces the number of input visual tokens.
| Language Model | Vision Encoder | VL Projector | Tokens | GQA | SQA | VQA | POPE | MME | MMB | ||
CLIP-B-P16-S224 [97] | MLP [75] | 196 | 55.3 | 52.8 | 38.3 | 83.2 | 1079.9 | 46.1 | ||
| CLIP-L-P14-S224 [97] | MLP [75] | 256 | 55.8 | 53.4 | 40.9 | 83.9 | 1104.4 | 50.9 | |||
| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 57.5 | 54.2 | 38.0 | 84.9 | 1123.3 | 53.9 | |||
| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 56.9 | 53.6 | 43.7 | 85.7 | 1137.7 | 52.8 | |||
| CLIP-L-P14-S336 [97] | LDP (ours) | 144 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 | |||
CLIP-B-P16-S224 [97] | MLP [75] | 196 | 57.0 | 58.4 | 43.1 | 83.8 | 1212.2 | 54.6 | ||
| CLIP-L-P16-S224 [97] | MLP [75] | 256 | 57.9 | 58.6 | 45.3 | 85.1 | 1285.0 | 57.7 | |||
| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 59.5 | 58.9 | 43.9 | 85.1 | 1275.6 | 59.5 | |||
| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 59.1 | 58.3 | 47.3 | 85.8 | 1333.1 | 57.1 | |||
| CLIP-L-P14-S336 [97] | LDP (ours) | 144 | 59.0 | 61.0 | 47.5 | 84.9 | 1288.9 | 59.6 |
| Vision Encoder | Method | Pretrain Paradigm | Pretrain Data | Tokens | GQA | SQA | VQA | POPE | MME | MMB |
| ViT-B-P16-S224 | SAIM [96] | UnSupervised | IN1K [33] | 196 | 45.3 | 40.9 | 32.2 | 76.5 | 845.4 | 2.1 |
| ViT-B-P16-S224 | MAE [47] | UnSupervised | IN1K [33] | 196 | 50.3 | 49.1 | 33.4 | 80.2 | 931.2 | 24.7 |
| ViT-B-P16-S224 | ViT [34] | Classification | IN21K [100] | 196 | 48.3 | 50.7 | 33.1 | 80.3 | 892.8 | 34.9 |
| Twins-SVT-L-S224 | Twins [22] | Classification | IN1K [33] | 49 | 50.3 | 50.9 | 33.2 | 80.6 | 941.6 | 32.1 |
| Twins-SVT-L-S384 | Twins [22] | Classification | IN1K [33] | 144 | 51.4 | 51.0 | 32.8 | 81.7 | 930.3 | 33.4 |
| Swin-Tiny-S224 | Swin [81] | Classification | IN1K [33] | 49 | 48.3 | 50.3 | 32.5 | 80.8 | 929.0 | 31.4 |
| Swin-Base-S384 | Swin [81] | Classification | IN22K [33] | 144 | 53.3 | 52.2 | 33.5 | 82.8 | 1037.2 | 40.3 |
| Swin-Tiny-S224 | GDino [79] | Grounding-Det | OGC [79] | 49 | 51.2 | 50.5 | 32.4 | 81.7 | 932.4 | 31.8 |
| Swin-Base-S384 | GDino [79] | Grounding-Det | COGOOR [79] | 144 | 54.9 | 51.0 | 33.8 | 84.5 | 1072.3 | 40.0 |
| ViT-B-P14-S224 | CLIP [97] | Image-Text-Align | WIT [97] | 256 | 55.3 | 52.8 | 38.3 | 83.2 | 1079.9 | 46.1 |
5 Ablation Study
5.1 Ablation on Vision Backbones
In Table 7, we compare the multimodal performance at varying model scales and different numbers of visual tokens. All experiments are conducted with CLIP ViT as a visual encoder. We configure different model scales, patch sizes, and types of vision-language projectors.
The impact of model scales. As the model scales up, the multimodal performance on 6 benchmarks maintains a gradual increase trend under the same projector [75]. However, it can be observed that the gain brought by the visual model scaling may gradually become saturated at a certain amount of training data.
The impact of the number of visual tokens. Compared with rows 4-5, our proposed LDP module reduces the number of visual tokens from 576 to 144 ( 75%), and it finally achieves performance equivalent to or sometimes better than the baseline. This reveals that the quality of visual tokens can be further improved while our proposed LDP module is quite effective.
The impact of pre-training paradigms. Furthermore, we show the performance of MobileLLaMA 1.4B under different vision backbone pre-training paradigms in Table 8. Based on the cost of annotation and pre-training, we roughly classify these paradigms into four categories. It turns out that the performance of MobileVLM gradually improves as the pre-training cost increases. The vision encoder pre-trained with supervised image-text alignment achieved the best performance. By comparing Swin-Base-S384-GDino and ViT-B-P14-S224, we notice that the model pre-trained by grounding detection achieved relatively comparable performance to the CLIP pre-trained model on GQA, SQA, POPE, and MME. This outcome indicates that the image-level alignment has greater potential to strike better performance than object-level, especially by using more visual tokens or more training data. In addition, better ImageNet performance of pre-trained models (e.g., Swin ViT) often corresponds to more general visual feature extraction capabilities, and MobileVLM will have certain performance gains in turn.
| VL Projector Architecture Design | Tokens | GQA | SQA | VQA | POPE | MME | MMB |
| 576 | 56.9 | 53.6 | 43.7 | 85.7 | 1137.7 | 52.8 | |
| 144 | 54.9 | 52.9 | 40.2 | 84.0 | 1150.8 | 50.3 | |
| 144 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 | |
| 144 | 55.3 | 53.9 | 40.8 | 84.6 | 1166.3 | 53.0 | |
| 144 | 55.6 | 54.3 | 41.5 | 84.6 | 1166.2 | 52.8 |
| LLM base model | SFT strategy | Conversation mode | GQA | SQA | VQA | POPE | MME | MMB |
| MobileLLaMA 1.4B | w/o | llava | 55.8 | 52.6 | 40.5 | 84.4 | 1111.5 | 52.0 |
| MobileLLaMA 1.4B | w/o | vicunav1 | 56.1 | 53.0 | 40.5 | 84.6 | 1148.5 | 50.3 |
| MobileLLaMA 1.4B | Alpaca | llava | 55.2 | 54.8 | 40.6 | 84.9 | 1171.1 | 51.9 |
| MobileLLaMA 1.4B | Alpaca | vicunav1 | 55.5 | 53.1 | 40.6 | 83.8 | 1168.0 | 47.7 |
| MobileLLaMA 1.4B | Vicuna | vicunav1 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 |
5.2 Abaltion on VL Projectors
Motivated by the fact both feature interaction and token interaction are beneficial, we utilize depthwise convolutions for the former and pointwise for the latter. Table 9 shows the performance of various VL projectors. Row in Table 9 is the module used in LLaVA [78] where only the feature space is transformed through two layers. Row adds a (depthwise) convolution before each (pointwise) for token interaction, which performs downsampling with a stride of . We notice that the performance begins to show an evident decline. Based on the setup of tokens, adding two front-end layers brings more feature-level interactions, which makes up for the performance loss caused by token reduction. Rows and show that adding more parameters does not achieve desired gains. Rows and show that the downsampling of tokens at the end of the projector has a positive effect.
5.3 Visual Resolution and Token Numbers
Since the number of visual tokens directly affects the inference speed of the whole multimodal model, we compare two types of design: reducing the input resolution (RIR) and using a lightweight downsample projector (LDP). Without loss of generality, for an image of with a patch size of , the former strategy generates tokens. For the latter, it produces tokens using a downsampling ratio of 2. We use for LDP and for RIR to keep the total number of tokens as 144. The result from Table 11 verifies the effectiveness of the proposed LDP.
| Method | GQA | SQA | VQA | POPE | MME | MMB |
| LDP | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 |
| RIR | 53.9 | 53.1 | 37.1 | 81.5 | 1072.5 | 46.7 |
5.4 Quantitative Analysis on SFT
Vicuna [132], fine-tuned on LLaMA, has been widely chosen in large multimodal models [75, 30, 133]. We further explore how SFT affects our language model’s performance in downstream tasks. Two common SFT paradigms Alpaca [111] and Vicuna [132] are compared in Table 10. We find that the scores of SQA, VQA, MME, and MMBench can all be significantly enhanced. It demonstrates that fine-tuning large language models in Vicuna dialogue mode [132] with the data from ShareGPT ultimately achieves the best performance. To better integrate SFT’s prompt format with the training of downstream tasks, we ablate the conversation mode on MobileVLM to find vicunav1 performs best.
6 Conclusion
In a nutshell, we present MobileVLM, a set of efficient and high-powered mobile scale vision language models tailored for mobile and IoT devices. In its making, we refurbish both language models and vision projection modules. Extensive experiments are conducted to choose proper vision backbones, to design an efficient projector, and to enhance model capacity by training schemes like language model SFT, a two-stage training strategy involving pretraining and instruction tuning, and LoRA fine-tuning. The performance is evaluated vigorously on mainstream VLM benchmarks. MobileVLMs also show an unprecedented speed on typical mobile and IoT devices. We believe that MobileVLM will open up new possibilities for widespread applications like multimodal assistants deployed on mobile devices or in self-driving cars, and more broadly embodied AI robots.
Acknowledgements:
This work was in part supported by National Key R&D Program of China (No. 2022ZD0118700).
The authors thank Yerui Sun for infrastructure support.
Appendix A Dataset
A.1 Dataset for MobileLLaMA
Table 12 shows the sampling strategy for the training set of MobileLLaMA.
| Dataset | Proportion | |
| Arxiv | 2.5% | |
| Book | 4.5% | |
| C4 | 15.0% | |
| Common Crawl | 67.7% | |
| Github | 4.5% | |
| StackExchange | 2.0% | |
| Wikipedia | 4.5% |
Appendix B SFT prompt and response example
We present our prompt template of MobileLLaMA SFT in Table 14.
| Human: | who are you? |
| MobileLLaMA 1.4B Base: | I’m the AI. |
| MobileLLaMA 1.4B SFT: | I am a digital asset management and metadata expert. My expertise includes creating and managing digital content, metadata, and digital asset management solutions. I have a passion for helping organizations unlock the value of their digital assets and improve their workflows. I am also experienced in working with big data and analytics and can help organizations unlock insights from their data to drive better decision-making. I am always up for a challenge and love working with clients to help them achieve their goals. |
| Human: | what can you do? |
| MobileLLaMA 1.4B Base: | I’m a chatbot. |
| MobileLLaMA 1.4B SFT: | In addition to providing expert advice and guidance on a wide range of topics, I can also help you with tasks and projects that are more complex or require specialized knowledge, such as developing and implementing project plans or managing budget and resource allocation. Overall, my role is to provide value and support to your project or task and to help you achieve your goals in a timely and efficient manner. |
| System | A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. |
| USER | I want you to act as my travel agent for preparing an itinerary for travel to Munnar and Tekkady in Kerala. I have already booked flights from Hyderabad to Kochi for an onward journey on 25th Jan and a return journey on 28th Jan. We are a group of 4 men and planning to stay 2 days in Munnar and 1 day in Tekkady. I want you to help me with accommodation booking and activities. |
[…]
Appendix C VLM demos
Figure 6 and Figure 7 showcase MobileVLM’s performance on each benchmark, including the scenes on attribute understanding (GQA left), spatial and relational reasoning (GQA right, MMBench), social science (ScienceQA left), natural science (ScienceQA right), OCR (TextVQA), object recognition (POPE), word knowledge (MME right), etc.

### v2

@@
 2.3 VLMs
-| Model | Vision Encoder | Language Model | Cross-modality Design | Multimodal Training Corpora |
-| CLIP [97] | ViT, ResNet | Transformer | Linear-Projection | WebImageText [97] (400M image-text pairs) |
-| BLIP [67] | ViT | MED∗ | Cross-Attention | COCO [73], VG [62], CC3M [103], CC12M [13], LAION [102] |
-| BLIP-2 [66] | CLIP/Eva-CLIP ViT | OPT, Flan-T5 | Q-Former | same as BLIP |
-| InstructBLIP [30] | ViT-G/14 | Flan-T5, Vicuna | Q-Former w/ FC | 13 held-in out of 26 public datasets |
-| Flamingo [2] | NFNet-F6 [59] | Chinchilla | Perceiver-Resampler [57] | M3W(43M), ALIGN(1.4B) [58], LTIP (312M), VTP (27M) |
-| LLaVA [76] | CLIP ViT-L/14 | Vicuna 7B/13B | Linear-Pprojection | filtered CC-595K from CC3M [103], LLaVA-Instruct-158K |
-| LLaVA-1.5 [74] | CLIP ViT-L@336 | Vicuna-7B/13B | MLP | a subset of InstructBLIP (1.2M) |
-| MiniGPT-4 [133] | EVA-CLIP ViT-G/14 | Vicuna-7B | Q-Former | LAION, CC, SBU [91] |
-| Shikra [15] | CLIP ViT-L/14 | Vicuna-7/13B | FC-layer | Flickr30K [95], RefCOCO [61], VG, Visual-7W [85] |
-| mPLUG-Owl [126] | CLIP ViT-L/14 | LLaMA-7B | Cross-Attention | LAION-400M, COYO-700M [11], CC, COCO |
-| KOSMOS-2 [94] | VLMo [6] | MAGNETO [119] | Perceiver-Resampler | GRIT (curated with LAION-2B, COYO-700M) |
-| QWen-VL [5] | Openclip ViT-bigG [54] | Qwen-LM | Cross-Attention | multi-tasking datasets (Captioning, VQA, Grounding, etc. ) |
-| ShareGPT4V [16] | CLIP ViT-L/14@336 | Vicuna-7B | MLP | ShareGPT4V (100K by GPT-4V, 1.2M by its learned model) |
-| MobileVLM (ours) | CLIP ViT-L/14@336 | MobileLLaMA | LDP | same as LLaVA-1.5 [74] |
+| Model | Vision Encoder | Language Model |
+Cross-modality Design |
+Multimodal Training Corpora |
+| CLIP [97] | ViT, ResNet | Transformer |
+Linear-Projection |
+WebImageText [97] (400M image-text pairs) |
+| BLIP [67] | ViT | MED |
+Cross-Attention |
+| BLIP-2 [66] | CLIP/Eva-CLIP ViT | OPT, Flan-T5 |
+Q-Former |
+same as BLIP |
+| InstructBLIP [30] | ViT-G/14 | Flan-T5, Vicuna |
+Q-Former w/ FC |
+13 held-in out of 26 public datasets |
+| Flamingo [2] | NFNet-F6 [59] | Chinchilla |
+Perceiver-Resampler [57] |
+M3W(43M), ALIGN(1.4B) [58], LTIP (312M), VTP (27M) |
+| LLaVA [76] | CLIP ViT-L/14 | Vicuna 7B/13B |
+Linear-Projection |
+filtered CC-595K from CC3M [103], LLaVA-Instruct-158K |
+| LLaVA-1.5 [74] | CLIP ViT-L@336 | Vicuna-7B/13B |
+MLP |
+a subset of InstructBLIP (1.2M) |
+| MiniGPT-4 [133] | EVA-CLIP ViT-G/14 | Vicuna-7B |
+Q-Former |
+LAION, CC, SBU [91] |
+| Shikra [15] | CLIP ViT-L/14 | Vicuna-7/13B |
+FC-layer |
+| mPLUG-Owl [126] | CLIP ViT-L/14 | LLaMA-7B |
+Cross-Attention |
+LAION-400M, COYO-700M [11], CC, COCO |
+| KOSMOS-2 [94] | VLMo [6] | MAGNETO [119] |
+Perceiver-Resampler |
+GRIT (curated with LAION-2B, COYO-700M) |
+| QWen-VL [5] | Openclip ViT-bigG [54] | Qwen-LM |
+Cross-Attention |
+multi-tasking datasets (Captioning, VQA, Grounding, etc. ) |
+| ShareGPT4V [16] | CLIP ViT-L/14@336 | Vicuna-7B |
+MLP |
+ShareGPT4V (100K by GPT-4V, 1.2M by its learned model) |
+| MobileVLM (ours) | CLIP ViT-L/14@336 | MobileLLaMA |
+LDP |
+same as LLaVA-1.5 [74] |
 Throughout recent years, a school of vision language models has rapidly emerged. Table 1 summarizes them in a detailed comparison regarding architectures, cross-modality design, and training corpora.
@@
 The construction of training data has become increasingly crucial.
-It is quite common to utilize millions of text-image pairs in the line of VLMs, where the new datasets are usually released alongside their corresponding new models.
+It is common to utilize millions of text-image pairs in the line of VLMs, where the new datasets are usually released alongside their corresponding new models.
 To name a few, apart from an enhanced visual receptor and novel language model called Qwen-LM [4], the multilingual multimodal Qwen-VL [5] additionally aligns the image with caption and box tuples, which sets a new record of generalist models.
@@
 In contrast, GQA [52] mainly centers around the VLMs’ abilities in real-world reasoning, scene understanding, and compositional question answering.
-TextVQA [105] consists of questions related to text in the image, evaluating the OCR and reasoning abilities of models.
+TextVQA [105] consists of questions related to the text in the image, evaluating the OCR and reasoning abilities of models.
 ScienceQA [83] consists of multimodal multiple-choice questions covering scientific topics, e.g., natural science, social science, and language science, which requires VLMs to integrate common sense into reasoning.
@@
 Formally, LDP (denoted as ) takes the visual embeddings as input, and outputs the efficiently extracted and aligned visual tokens as,
-| Model | Common Sense Reasoning | Language Understanding | ||||||||
-| ARC(c/e) | BoolQ | RTE | Winogrande | TruthfulQA | HellaSwag | PIQA | MMLU | Avg. | ||
-| INCITE 3B (V1) [29] | 0.32 / 0.68 | 0.67 | 0.52 | 0.63 | 0.21 / 0.33 | 0.48 | 0.74 | 0.2675 | 0.4848 | |
-| OpenLLaMA 3B (V1) [43] | 0.34 / 0.69 | 0.68 | 0.58 | 0.62 | 0.22 / 0.35 | 0.49 | 0.75 | 0.2695 | 0.4990 | |
-| MobileLLaMA 2.7B | 0.32 / 0.68 | 0.61 | 0.59 | 0.63 | 0.23 / 0.36 | 0.48 | 0.75 | 0.2730 | 0.4923 | |
-| TinyLLaMA 1.1B (2T) [93] | 0.23 / 0.57 | 0.59 | 0.55 | 0.57 | 0.23 / 0.39 | 0.40 | 0.70 | 0.2541 | 0.4484 | |
-| Galactica 1.3B [112] | 0.28 / 0.62 | 0.62 | 0.52 | 0.55 | 0.25 / 0.41 | 0.34 | 0.63 | 0.2675 | 0.4488 | |
-| OPT 1.3B [131] | 0.24 / 0.57 | 0.56 | 0.51 | 0.59 | 0.24 / 0.39 | 0.41 | 0.71 | 0.2461 | 0.4466 | |
-| Pythia 1.4B [7] | 0.26 / 0.61 | 0.63 | 0.52 | 0.57 | 0.23 / 0.39 | 0.40 | 0.71 | 0.2568 | 0.4577 | |
-| MobileLLaMA 1.4B | 0.26 / 0.61 | 0.53 | 0.53 | 0.59 | 0.21 / 0.35 | 0.43 | 0.73 | 0.2497 | 0.4490 | |
+| Model | Common Sense Reasoning | Language Understanding | |||||||
+| ARC | BoolQ | RTE | Winogrande | TruthfulQA | HellaSwag | PIQA | MMLU | Avg. | |
+| INCITE 3B (V1) [29] | 0.32 / 0.68 | 0.67 | 0.52 | 0.63 | 0.21 / 0.33 | 0.48 | 0.74 | 0.2675 | 0.4848 |
+| OpenLLaMA 3B (V1) [43] | 0.34 / 0.69 | 0.68 | 0.58 | 0.62 | 0.22 / 0.35 | 0.49 | 0.75 | 0.2695 | 0.4990 |
+| MobileLLaMA 2.7B | 0.32 / 0.68 | 0.61 | 0.59 | 0.63 | 0.23 / 0.36 | 0.48 | 0.75 | 0.2730 | 0.4923 |
+| TinyLLaMA 1.1B (2T) [93] | 0.23 / 0.57 | 0.59 | 0.55 | 0.57 | 0.23 / 0.39 | 0.40 | 0.70 | 0.2541 | 0.4484 |
+| Galactica 1.3B [112] | 0.28 / 0.62 | 0.62 | 0.52 | 0.55 | 0.25 / 0.41 | 0.34 | 0.63 | 0.2675 | 0.4488 |
+| OPT 1.3B [131] | 0.24 / 0.57 | 0.56 | 0.51 | 0.59 | 0.24 / 0.39 | 0.41 | 0.71 | 0.2461 | 0.4466 |
+| Pythia 1.4B [7] | 0.26 / 0.61 | 0.63 | 0.52 | 0.57 | 0.23 / 0.39 | 0.40 | 0.71 | 0.2568 | 0.4577 |
+| MobileLLaMA 1.4B | 0.26 / 0.61 | 0.53 | 0.53 | 0.59 | 0.21 / 0.35 | 0.43 | 0.73 | 0.2497 | 0.4490 |
 4 Experiment
@@
 The whole reproducible training process is composed of three stages.
-Firstly, we pretrain LLM foundation models on the text-only dataset RedPajama v1 [29].
+Firstly, we pre-train LLM foundation models on the text-only dataset RedPajama v1 [29].
 Secondly, we perform supervised fine-tuning (SFT) following Vicuna [20] on a dataset of multi-turn dialogues between humans and ChatGPT from third-party platforms.
@@
 We utilize a global batch size of 5,242,880.
-The peak learning rate is set to 3 10-4 and it decays to 3 10-5 following the cosine strategy.
+The peak learning rate is set to 3 10 and it decays to 3 10 following the cosine strategy.
 We warm up with 2000 iterations.
@@
 At the 3B level, our MobileLLaMA 2.7B also demonstrates competitive performance to INCITE 3B (V1) and OpenLLaMA 3B (V1), while MobileLLaMA 2.7B being about 40% faster than OpenLLaMA 3B on a Snapdragon 888 CPU as shown in Table 5.
-For common sense reasoning, we report the zero-shot accuracy on five prevalent benchmarks, i.e., ARCe [10], ARCc [28], BoolQ [27], Winogrande [101], and TruthfulQA [72].
+For common sense reasoning, we report the zero-shot accuracy on five prevalent benchmarks, i.e., ARC [10], ARC [28], BoolQ [27], Winogrande [101], and TruthfulQA [72].
 Our models demonstrate strong reasoning capabilities on these benchmarks.
@@
 The superior language understanding capability makes our models more suitable for downstream tasks, particularly for instruction tuning, in-context learning, etc.
-| Model | Hardware | Precision |
-| OpenLLaMA 3B | Snapdragon 888 | 8-bit | 3.4 | 3093 | 7.32 | 6.58 | 63.33 | ||||||||
-| Snapdragon 888 | 4-bit | 2.3 | 3604 | 8.97 | 7.14 | 58.04 | |||||||||
-| MobileLLaMA 2.7B | Snapdragon 888 | 8-bit | 2.7 | 3919 | 17.59 | 9.14 | 44.85 | ||||||||
-| Snapdragon 888 | 4-bit | 1.5 | 3932 | 18.10 | 14.71 | 28.30 | |||||||||
-| TinyLLaMA 1.1B | Snapdragon 888 | 8-bit | 1.1 | 4215 | 39.49 | 19.75 | 20.83 | ||||||||
-| Snapdragon 888 | 4-bit | 0.6 | 3887 | 44.17 | 31.54 | 13.22 | |||||||||
-| MobileLLaMA 1.4B | Snapdragon 888 | 8-bit | 1.4 | 3846 | 35.46 | 17.93 | 22.81 | ||||||||
-| Snapdragon 888 | 4-bit | 0.7 | 3870 | 36.20 | 28.32 | 14.76 | |||||||||
-| OpenLLaMA 3B | Jetson Orin | 8-bit | 3.4 | 2382 | 80.34 | 29.97 | 13.94 | ||||||||
-| Jetson Orin | 4-bit | 2.3 | 3340 | 143.25 | 32.16 | 12.83 | |||||||||
-| MobileLLaMA 2.7B | Jetson Orin | 8-bit | 2.7 | 3040 | 133.41 | 33.28 | 12.46 | ||||||||
-| Jetson Orin | 4-bit | 1.5 | 2647 | 130.97 | 38.99 | 10.74 | |||||||||
-| TinyLLaMA 1.1B | Jetson Orin | 8-bit | 1.1 | 3007 | 279.61 | 72.30 | 5.89 | ||||||||
-| Jetson Orin | 4-bit | 0.6 | 3801 | 306.76 | 78.83 | 5.38 | |||||||||
-| MobileLLaMA 1.4B | Jetson Orin | 8-bit | 1.4 | 3289 | 249.56 | 60.73 | 6.96 | ||||||||
+| Model | Hardware | Precision | Size (GB) | (tokens/s) | (tokens/s) | (tokens/s) | (s) |
+| OpenLLaMA 3B | Snapdragon 888 | 8-bit | 3.4 | 3093 | 7.32 | 6.58 | 63.33 |
+| Snapdragon 888 | 4-bit | 2.3 | 3604 | 8.97 | 7.14 | 58.04 | |
+| MobileLLaMA 2.7B | Snapdragon 888 | 8-bit | 2.7 | 3919 | 17.59 | 9.14 | 44.85 |
+| Snapdragon 888 | 4-bit | 1.5 | 3932 | 18.10 | 14.71 | 28.30 | |
+| TinyLLaMA 1.1B | Snapdragon 888 | 8-bit | 1.1 | 4215 | 39.49 | 19.75 | 20.83 |
+| Snapdragon 888 | 4-bit | 0.6 | 3887 | 44.17 | 31.54 | 13.22 | |
+| MobileLLaMA 1.4B | Snapdragon 888 | 8-bit | 1.4 | 3846 | 35.46 | 17.93 | 22.81 |
+| Snapdragon 888 | 4-bit | 0.7 | 3870 | 36.20 | 28.32 | 14.76 | |
+| OpenLLaMA 3B | Jetson Orin | 8-bit | 3.4 | 2382 | 80.34 | 29.97 | 13.94 |
+| Jetson Orin | 4-bit | 2.3 | 3340 | 143.25 | 32.16 | 12.83 | |
+| MobileLLaMA 2.7B | Jetson Orin | 8-bit | 2.7 | 3040 | 133.41 | 33.28 | 12.46 |
+| Jetson Orin | 4-bit | 1.5 | 2647 | 130.97 | 38.99 | 10.74 | |
+| TinyLLaMA 1.1B | Jetson Orin | 8-bit | 1.1 | 3007 | 279.61 | 72.30 | 5.89 |
+| Jetson Orin | 4-bit | 0.6 | 3801 | 306.76 | 78.83 | 5.38 | |
+| MobileLLaMA 1.4B | Jetson Orin | 8-bit | 1.4 | 3289 | 249.56 | 60.73 | 6.96 |
 | Jetson Orin | 4-bit | 0.7 | 3738 | 253.22 | 66.79 | 6.33 |
@@
 We can see that MobileVLM with LoRA achieves comparable performances to those of full finetuning on 6 benchmarks, which is consistent with [51].
-| Model | LM | Hardware |
-Vicuna 7B | Snapdragon | 4.70 | 8.23 | 17347 | 5.36 | 0.25 | 329.89 | ||||||||||
-| OpenLLaMA 3B | Snapdragon | 2.88 | 7.98 | 27530 | 8.95 | 7.22 | 84.43 | |||||||||||
-| TinyLLaMA 1B | Snapdragon | 1.18 | 7.77 | 31370 | 41.70 | 18.40 | 20.70 | |||||||||||
-MobileLLaMA 2.7B | Snapdragon | 2.14 | 8.43 | 27660 | 18.36 | 12.21 | 33.10 | ||||||||||
-| MobileLLaMA 1.4B | Snapdragon | 1.40 | 6.82 | 34892 | 34.93 | 21.54 | 18.51 | |||||||||||
-Vicuna 7B | Jetson Orin | 4.70 | 2.89 | 9281 | 367.26 | 17.74 | 19.75 | ||||||||||
-| OpenLLaMA 3B | Jetson Orin | 2.88 | 2.94 | 22270 | 474.49 | 30.66 | 12.52 | |||||||||||
-| TinyLLaMA 1B | Jetson Orin | 1.18 | 2.98 | 24655 | 1253.94 | 76.63 | 5.90 | |||||||||||
-MobileLLaMA 2.7B | Jetson Orin | 2.14 | 3.11 | 15678 | 440.60 | 38.34 | 8.31 | ||||||||||
+| Model | LM | Hardware | Size (GB) | (ms/patch) | (tokens/s) | (tokens/s) | (tokens/s) | (s) |
+| LLaVA-v1.5-336 | Vicuna 7B | Snapdragon | 4.70 | 8.23 | 17347 | 5.36 | 0.25 | 329.89 |
+| OpenLLaMA 3B | Snapdragon | 2.88 | 7.98 | 27530 | 8.95 | 7.22 | 84.43 | |
+| TinyLLaMA 1B | Snapdragon | 1.18 | 7.77 | 31370 | 41.70 | 18.40 | 20.70 | |
+| MobileVLM-336 | MobileLLaMA 2.7B | Snapdragon | 2.14 | 8.43 | 27660 | 18.36 | 12.21 | 33.10 |
+| MobileLLaMA 1.4B | Snapdragon | 1.40 | 6.82 | 34892 | 34.93 | 21.54 | 18.51 | |
+| LLaVA-v1.5-336 | Vicuna 7B | Jetson Orin | 4.70 | 2.89 | 9281 | 367.26 | 17.74 | 19.75 |
+| OpenLLaMA 3B | Jetson Orin | 2.88 | 2.94 | 22270 | 474.49 | 30.66 | 12.52 | |
+| TinyLLaMA 1B | Jetson Orin | 1.18 | 2.98 | 24655 | 1253.94 | 76.63 | 5.90 | |
+| MobileVLM-336 | MobileLLaMA 2.7B | Jetson Orin | 2.14 | 3.11 | 15678 | 440.60 | 38.34 | 8.31 |
 | MobileLLaMA 1.4B | Jetson Orin | 1.40 | 3.32 | 17712 | 667.69 | 65.27 | 5.14 |
@@
 This precisely proves the importance of efficient projection that reduces the number of input visual tokens.
-| Language Model | Vision Encoder | VL Projector | Tokens | GQA | SQA | VQA | POPE | MME | MMB | ||
-CLIP-B-P16-S224 [97] | MLP [75] | 196 | 55.3 | 52.8 | 38.3 | 83.2 | 1079.9 | 46.1 | ||
-| CLIP-L-P14-S224 [97] | MLP [75] | 256 | 55.8 | 53.4 | 40.9 | 83.9 | 1104.4 | 50.9 | |||
-| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 57.5 | 54.2 | 38.0 | 84.9 | 1123.3 | 53.9 | |||
-| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 56.9 | 53.6 | 43.7 | 85.7 | 1137.7 | 52.8 | |||
-| CLIP-L-P14-S336 [97] | LDP (ours) | 144 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 | |||
-CLIP-B-P16-S224 [97] | MLP [75] | 196 | 57.0 | 58.4 | 43.1 | 83.8 | 1212.2 | 54.6 | ||
-| CLIP-L-P16-S224 [97] | MLP [75] | 256 | 57.9 | 58.6 | 45.3 | 85.1 | 1285.0 | 57.7 | |||
-| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 59.5 | 58.9 | 43.9 | 85.1 | 1275.6 | 59.5 | |||
-| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 59.1 | 58.3 | 47.3 | 85.8 | 1333.1 | 57.1 | |||
+| Language Model | Vision Encoder | VL Projector | Tokens | GQA | SQA | VQA | POPE | MME | MMB |
+| MobileLLaMA 1.4B | CLIP-B-P16-S224 [97] | MLP [75] | 196 | 55.3 | 52.8 | 38.3 | 83.2 | 1079.9 | 46.1 |
+| CLIP-L-P14-S224 [97] | MLP [75] | 256 | 55.8 | 53.4 | 40.9 | 83.9 | 1104.4 | 50.9 | |
+| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 57.5 | 54.2 | 38.0 | 84.9 | 1123.3 | 53.9 | |
+| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 56.9 | 53.6 | 43.7 | 85.7 | 1137.7 | 52.8 | |
+| CLIP-L-P14-S336 [97] | LDP (ours) | 144 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 | |
+| MobileLLaMA 2.7B | CLIP-B-P16-S224 [97] | MLP [75] | 196 | 57.0 | 58.4 | 43.1 | 83.8 | 1212.2 | 54.6 |
+| CLIP-L-P16-S224 [97] | MLP [75] | 256 | 57.9 | 58.6 | 45.3 | 85.1 | 1285.0 | 57.7 | |
+| CLIP-G-P14-S224 [110] | MLP [75] | 256 | 59.5 | 58.9 | 43.9 | 85.1 | 1275.6 | 59.5 | |
+| CLIP-L-P14-S336 [97] | MLP [75] | 576 | 59.1 | 58.3 | 47.3 | 85.8 | 1333.1 | 57.1 | |
 | CLIP-L-P14-S336 [97] | LDP (ours) | 144 | 59.0 | 61.0 | 47.5 | 84.9 | 1288.9 | 59.6 |
@@
 | MobileLLaMA 1.4B | w/o | llava | 55.8 | 52.6 | 40.5 | 84.4 | 1111.5 | 52.0 |
-| MobileLLaMA 1.4B | w/o | vicunav1 | 56.1 | 53.0 | 40.5 | 84.6 | 1148.5 | 50.3 |
+| MobileLLaMA 1.4B | w/o | vicuna | 56.1 | 53.0 | 40.5 | 84.6 | 1148.5 | 50.3 |
 | MobileLLaMA 1.4B | Alpaca | llava | 55.2 | 54.8 | 40.6 | 84.9 | 1171.1 | 51.9 |
-| MobileLLaMA 1.4B | Alpaca | vicunav1 | 55.5 | 53.1 | 40.6 | 83.8 | 1168.0 | 47.7 |
-| MobileLLaMA 1.4B | Vicuna | vicunav1 | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 |
+| MobileLLaMA 1.4B | Alpaca | vicuna | 55.5 | 53.1 | 40.6 | 83.8 | 1168.0 | 47.7 |
+| MobileLLaMA 1.4B | Vicuna | vicuna | 56.1 | 54.7 | 41.5 | 84.5 | 1196.2 | 53.2 |
 5.2 Abaltion on VL Projectors
@@
 It demonstrates that fine-tuning large language models in Vicuna dialogue mode [132] with the data from ShareGPT ultimately achieves the best performance.
-To better integrate SFT’s prompt format with the training of downstream tasks, we ablate the conversation mode on MobileVLM to find vicunav1 performs best.
+To better integrate SFT’s prompt format with the training of downstream tasks, we ablate the conversation mode on MobileVLM to find vicuna performs best.
 6 Conclusion
@@
 Table 12 shows the sampling strategy for the training set of MobileLLaMA.
-| Dataset | Proportion | |
-| Arxiv | 2.5% | |
-| Book | 4.5% | |
-| C4 | 15.0% | |
-| Common Crawl | 67.7% | |
-| Github | 4.5% | |
-| StackExchange | 2.0% | |
+| Dataset | Proportion |
+| Arxiv | 2.5% |
+| Book | 4.5% |
+| C4 | 15.0% |
+| Common Crawl | 67.7% |
+| Github | 4.5% |
+| StackExchange | 2.0% |
 | Wikipedia | 4.5% |
@@
 We present our prompt template of MobileLLaMA SFT in Table 14.
-| Human: | who are you? |
-| MobileLLaMA 1.4B Base: | I’m the AI. |
-| MobileLLaMA 1.4B SFT: | I am a digital asset management and metadata expert.
+Human: |
+who are you? |
+MobileLLaMA 1.4B Base: |
+I’m the AI. |
+MobileLLaMA 1.4B SFT: |
+I am a digital asset management and metadata expert.
 My expertise includes creating and managing digital content, metadata, and digital asset management solutions.
@@
 |
-| Human: | what can you do? |
-| MobileLLaMA 1.4B Base: | I’m a chatbot. |
-| MobileLLaMA 1.4B SFT: | In addition to providing expert advice and guidance on a wide range of topics, I can also help you with tasks and projects that are more complex or require specialized knowledge, such as developing and implementing project plans or managing budget and resource allocation.
+Human: |
+what can you do? |
+MobileLLaMA 1.4B Base: |
+I’m a chatbot. |
+MobileLLaMA 1.4B SFT: |
+In addition to providing expert advice and guidance on a wide range of topics, I can also help you with tasks and projects that are more complex or require specialized knowledge, such as developing and implementing project plans or managing budget and resource allocation.
 Overall, my role is to provide value and support to your project or task and to help you achieve your goals in a timely and efficient manner.
 |
-| System | A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. |
-| USER | I want you to act as my travel agent for preparing an itinerary for travel to Munnar and Tekkady in Kerala.
+System |
+A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions. |
+USER |
+I want you to act as my travel agent for preparing an itinerary for travel to Munnar and Tekkady in Kerala.
 I have already booked flights from Hyderabad to Kochi for an onward journey on 25th Jan and a return journey on 28th Jan.
@@
 |
+ASSISTANT |
 […]